# WOMD-Reasoning: A Large-Scale Language Dataset for Interaction and Driving Intentions Reasoning

### v1

@@ A.2 Example of Translated Motion Scenario
The following is the description of the ego and surrounding agents after 3.0 seconds: Surrounding agent # 0 will be 8 meters in front of the ego agent, and will be 5 meters on the left of the ego agent. It will be heading right of the ego agent. Its speed will be 7 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 18 meters away from the intersection center. The ego agent will be 13 meters in front of the current place, and will be 2 meters on the right of the current place. It will be heading in the same direction as the current moment. Its speed will be 3 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 20 meters away from the intersection center. Surrounding agent # 2 will be 13 meters on the right of the ego agent, and will be 9 meters behind the ego agent. It will be heading left of the ego agent. It will not be moving. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 34 meters away from the intersection center.
$\cdots$
@@
A.3 Prompts for GPT
A.3.1 System Prompt
Table  6 presents the system prompt part of our prompt set.
@@ A.3.1 System Prompt
A.3.2 Responsibility Prompt
Table  7 presents the responsibility prompt part of our prompt set.
@@ A.3.2 Responsibility Prompt
A.3.3 Rule Prompt
Table  8 presents the rule prompt part of our prompt set.
Global rules prompt for GPT:
Here are some global rules that your need to follow: 1. Please provide the summary and analysis of the driving scenario according to your responsibilities. 2. Never use bullet points. Be concise and compact. 3. Give natural language, no parenthesis. 4. Give the results in the order of the responsibilities. Always start a new paragraph for each responsibility, but one responsibility can take several paragraphs. Do not combine multiple responsibilities together. Finish one responsibility before starting another. 5. Do not show the intersection number, lane number or the ego agent’s number. The surrounding agents’ numbers need to be shown. 6. Output in full. Never omit anything regardless of any reason. You are allowed to give very long outputs. 7. Remember that all the agents and traffic controls in the scene have been provided, so if any agents or traffic controls are not mentioned in the scene, they do not exist. You don’t need to consider them or give any output about them. Therefore, do not output words like "not mentioned" because this is inaccurate. 8. The traffic light info is for the current moment only. The future traffic light info is unknown. 9. Any action caused by traffic controls like lights, stop signs or crosswalks may also constitute yielding interactions. Traffic controls are only an indication of right-of-way, but any actions that are caused by the traffic controls are still considered interactions. For example, if a car is waiting for the red light or a stop sign while someone other cross the intersection with green light, the car is still considered yielding to the other car. 10. When analyzing the future, talk about the period between the current and future moment provided, not the two moments.
@@ A.3.3 Rule Prompt
A.3.4 In-context Prompt
Table  9 presents the in-context prompt part of our prompt set.

### v2

@@ A.1.1 Automatic Translation of Driving Scenarios into Language Descriptions
Descriptions on map environment: The ego agent is heading towards intersection. The intersection center is 18.0 meters in front of the ego agent, and is 17.0 meters on the left of the ego agent. The intersection is a 4 way intersection. Descriptions on ego agent: The ego agent is on the 1 lane from the right, out of 3 lanes. Its current speed is 8 m/s. It is accelerating. Traffic Light for the ego agent is red. The ego agent is approaching a crosswalk 3 meters ahead. Descriptions on surrounding (other) agents: Surrounding agent # 0 is a vehicle. It is 23 meters in front of the ego agent, and is 17 meters on the left of the ego agent. It is heading right of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. It is approaching a crosswalk 3 meters ahead. Surrounding agent # 2 is a vehicle. It is 11 meters on the right of the ego agent, and is 0 meters behind the ego agent. It is heading left of the ego agent. It is not moving. It is on the same side of the intersection as the ego agent. It is 34 meters away from the intersection center. $\cdots$ Descriptions on agents after 3 seconds: The following is the description of the ego and surrounding agents after 3.0 seconds: Surrounding agent # 0 will be 8 meters in front of the ego agent, and will be 5 meters on the left of the ego agent. It will be heading right of the ego agent. Its speed will be 7 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 18 meters away from the intersection center. The ego agent will be 13 meters in front of the current place, and will be 2 meters on the right of the current place. It will be heading in the same direction as the current moment. Its speed will be 3 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 20 meters away from the intersection center. Surrounding agent # 2 will be 13 meters on the right of the ego agent, and will be 9 meters behind the ego agent. It will be heading left of the ego agent. It will not be moving. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 34 meters away from the intersection center. $\cdots$
@@
A.1.2 GPT-based Interaction Analysis
With the language translation of WOMD as input, we further design a set of prompts to utilize ChatGPT-4 [34] to build the Q&A dataset, taking advantage of its abilities in analyzing interactions. The prompts come with 4 main parts. (1) System Prompt, which lets the GPT know its role and the format of its input, i.e. the language translation. (2) Responsibility Prompt, which describes the responsibilities of the GPT, including the questions it should ask and answer, as well as the output format it should use. This will be talked about in detail in the next paragraph. (3) Rules Prompt, which contains global rules to guide the output and the analysis. (4) In-context Prompt, which provides a few human-written examples of input-output pairs to perform in-context learning. An example of each part of the prompts can be found in Table 9, 10, 11, and 12 respectively.
The questions in the responsibility part of the prompt decide what would be present in the generated language Q&A dataset. To provide a comprehensive language dataset, and to present interaction analysis as accurately as possible, we organize the questions into a chain to enable a well-designed Chain-of-Thought [55, 55, 8]: First, we ask questions about the map environment, the ego and other agents’ motion. These questions would let the GPT provide Q&As useful for fine-tuning scene description agents while enhancing the GPT’s understanding of the input information at the same time. Then, we ask questions about the interactions that occur in the observable future period. To include traffic rule-induced and intention-induced interactions, we ask the GPT to think about the interaction between each pair of agents by answering a sequence of questions: Q1: Are the two agents vehicles, and are they close to each other with traffic rules or signals governing their movements? If the answer is yes, we request that GPT analyze their yielding relations according to the traffic rules. Otherwise, we ask Q2: Are the two agents vehicle and pedestrian respectively with intention conflicts? If so, the vehicle generally should yield to the pedestrian for safety. If both questions do not help find the interaction relations, we come to Q3: Does the scene show patterns of certain intention-induced interaction? We provide descriptions of a few common interactions like overtaking, following, etc. GPT can choose the best fit from these patterns to try to cover more interaction types. Based on this pipeline, we can provide interaction analysis with rich information on traffic rule-induced and human intention-induced interactions. After that, we can finally wrap up the interactions between the ego and all other agents to provide the intention of the ego agent.
@@ A.1.2 GPT-based Interaction Analysis
Global rules prompt for GPT: Here are some global rules that your need to follow: 1. Please provide the summary and analysis of the driving scenario according to your responsibilities. 2. Never use bullet points. Be concise and compact. 3. Give natural language, no parenthesis. 4. Give the results in the order of the responsibilities. Always start a new paragraph for each responsibility, but one responsibility can take several paragraphs. Do not combine multiple responsibilities together. Finish one responsibility before starting another. 5. Do not show the intersection number, lane number or the ego agent’s number. The surrounding agents’ numbers need to be shown. 6. Output in full. Never omit anything regardless of any reason. You are allowed to give very long outputs. 7. Remember that all the agents and traffic controls in the scene have been provided, so if any agents or traffic controls are not mentioned in the scene, they do not exist. You don’t need to consider them or give any output about them. Therefore, do not output words like "not mentioned" because this is inaccurate. 8. The traffic light info is for the current moment only. The future traffic light info is unknown. 9. Any action caused by traffic controls like lights, stop signs or crosswalks may also constitute yielding interactions. Traffic controls are only an indication of right-of-way, but any actions that are caused by the traffic controls are still considered interactions. For example, if a car is waiting for the red light or a stop sign while someone other cross the intersection with green light, the car is still considered yielding to the other car. 10. When analyzing the future, talk about the period between the current and future moment provided, not the two moments.
@@
Figure 7: More visual examples of WOMD-Reasoning through simulations.
A.2 Examples of Simulated Vision Modality
Figure 7 provides some whole examples of the simulated scenarios generated by MetaDrive [24] simulator using ScenarioNet [25] traffic scenario simulation platform.
A.3 Prompts for evaluating GPT Score
We employ the OpenAI GPT-4-turbo model [34] to evaluate the performance of all models benchmarked on the WOMD-Reasoning. The GPT Score prompt is presented in Table 13. Notably, we design our GPT Score prompt using a similar template to the one from [7].
We would like to request your feedback on the performance of an AI assistant in response to the user question displayed above. The user asks the question on observing a driving scenario. The correct answer is included for your reference. ### Scoring rules: Your scores for each answer should be between 0 (worst) and 10 (best). If the answer is mostly incorrect but includes important relevant information, such as identifying correct agents or actions, give a score of less than 3. Give intermediate scores for partially correct answers or when the numbers are close. Only give a score of 10 for flawless answers. Think critically and carefully. Most importantly, check the answer for factual correctness, but reward partial credit where applicable. ### Grading process For each of the xxx questions, provide a one line assessment of the student’s answer in the format: “‘ n. <short one sentence assessment>. Score: <score 0-10>. “‘ A few examples of what a good assessment might look like: 1. Acknowledged question unrelated to driving and attempted to answer. Score: 10. 2. Doesn’t answer the question but has all information necessary. Score: 3. 3. Incorrectly stated there are no vehicles around even though there is one. Score: 0. 4. Unable to answer question given information available. Score: 10. 5. Give 12 m/s or m or percentage for the correct answer of 13 (within 10% deviation). Score: 9. 6. Give 15 m/s or m or percentage for the correct answer of 13 (within 20% deviation). Score: 7. 7. Give 18 m/s or m or percentage for the correct answer of 13 (within 40% deviation). Score: 5. 8. Give 2 m/s or m or percentage for the correct answer of 13. Score: 0.
Table 13: Prompt for GPT Score.
Table 14: Refinement of Motion-LLaVA training strategy.
Multi-modal Fine-tuning Strategy Projector Alignment Encoder Freezing ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) Acc1.0 ($\uparrow$) Median ($\downarrow$) GPT Score ($\uparrow$)
LLaVA 0.821 0.678 0.484 5.82 0.771 2.3% 12.00 6.17
$\bm{-}$ Projector Alignment 0.821 0.697 0.492 5.96 0.782 2.3% 8.06 6.49
$\bm{-}$ Encoder Freezing (Ours) 0.840 0.736 0.516 6.35 0.794 56.2% 1.00 7.09
A.4 Qualitative Comparison of Benchmarked Multi-modal Models on Answering Questions in WOMD-Reasoning
In this appendix, we present examples of answers to questions in WOMD-Reasoning generated by the benchmarked multi-modal models listed in Table 4. The selected scenario corresponds to Figure 5, which already showcases answers from Motion-LLaVA. Additionally, we provide results for both the Non-Fine-tuned (i.e., off-the-shelf) and Fine-tuned LLaVA models, both using BEV plots as vision input.
For the Non-Fine-tuned LLaVA, we perform prompt engineering to make the answers coherent. However, the low GPT scores indicate that it remains challenging for LLaVA to directly comprehend the scenario:
  • Q: Can you describe the type of intersection present in the current driving scenario?
    A: The type of intersection present in the current driving scenario is a four-way intersection.
    GPT Score: 0. Incorrectly identifies the type of intersection.
  • Q: What is the ego agent’s current motion status?
    A: The ego agent is currently stopped at the intersection.
    GPT Score: 0. Incorrectly describes the motion status of the ego agent.
  • Q: Where is surrounding agent #5 in relation to the ego agent?
    A: Surrounding agent #5 is located to the right of the ego agent.
    GPT Score: 0. Incorrect location of surrounding agent #5 in relation to the ego agent.
  • Q: What kind of interaction will occur between the ego agent and surrounding agent #5?
    A: The ego agent will pass by the surrounding agent #5 at the intersection.
    GPT Score: 0. Incorrect prediction of the interaction between the ego agent and surrounding agent #5.
Quantitatively, the Non-Fine-tuned LLaVA achieves a remarkably low GPT Score of 2.31 across all questions in the validation set, demonstrating difficulty in understanding the driving scenarios based solely on the prior knowledge of the Vision-Language Model (VLM). For comparison, we evaluate the performance of the Fine-tuned LLaVA with vision input under the same scenario. The fine-tuned model achieves much more reasonable results, as illustrated below:
  • Q: Can you describe the type of intersection present in the current driving scenario?
    A: The intersection is a 3 way intersection.
    GPT Score: 10. The student’s answer matches the correct answer exactly.
  • Q: What is the ego agent’s current motion status?
    A: The ego agent is accelerating.
    GPT Score: 10. The student’s answer matches the correct answer exactly.
  • Q: Where is surrounding agent #5 in relation to the ego agent?
    A: It is 14 meters on the right and 10 meters in front of the ego agent.
    GPT Score: 0. The student’s answer incorrectly describes the position of surrounding agent #5.
  • Q: What kind of interaction will occur between the ego agent and surrounding agent #5?
    A: Surrounding agent #5 will follow the ego agent as they are both heading towards the intersection and are on the same side.
    GPT Score: 5. The student’s answer is partially correct but does not specify that agent #5 is behind the ego agent.
The results indicate that fine-tuning significantly improves LLaVA’s ability to interpret driving scenarios. This improvement can be attributed to the inclusion of our WOMD-Reasoning dataset. However, despite these enhancements, notable under-performance persists in questions related to Other Agent Status and Interaction & Intention, especially when compared to the results in Figure 5, which are generated using Motion-LLaVA. We attribute this limitation to potential information loss during the conversion of vectorized motion data into a BEV plot with a scalar bar.
A.5 Example Aggregated Context for Chain-of-Thought Reasoning
We present an example of an aggregated context, derived from the answers generated by Motion-LLaVA for factual questions, which is used by Motion-LLaVA to infer answers to reasoning-based questions. This context corresponds to the CoT configuration in Table 7:
The ego agent is turning left. Its current speed is 5 m/s. It is accelerating.
Surrounding agent #7 is a vehicle. It is 26 meters on the left of the ego agent and 8 meters in front. It is heading right of the ego agent. Its current speed is 3 m/s. It is decelerating.
The blue and orange sections contain answers to questions appearing in WOMD-Reasoning: blue addresses the Road Environment and Ego Agent Status, while orange covers Other Agent Status.
To generate this context, we prompt the Motion-LLaVA model with relevant questions and then concatenate the answers to create a complete context.
A.6 Example Factual Description Context for Chain-of-Thought Reasoning
We provide an example of a factual description context, generated using our motion data translation program as part of constructing WOMD-Reasoning, as described in Appendix A.1.1. This context is used by Motion-LLaVA to infer answers to reasoning-based questions and corresponds to the CoT w/ Facts Contexts configuration in Table 7. Notably, these contexts exclude future information compared to the translated motion data, thereby avoiding any information leakage:
The ego agent is in intersection. The intersection center is 6.0 meters in front of the ego agent, and is 0.0 meters on the left of the ego agent. The intersection is a 4 way intersection. It is turning left. It is exiting the intersection. Its current speed is 3 m/s. It is accelerating. Traffic Light for the ego agent is green. The ego agent is approaching a crosswalk 3 meters ahead.
Surrounding agent #7 is a vehicle. It is 26 meters on the left of the ego agent, and is 8 meters in front of the ego agent. It is heading right of the ego agent. Its current speed is 4 m/s. It is decelerating. It is heading towards the intersection. It is on the left of the intersection. Traffic Light for it is red 14.5 meters ahead. It is 26 meters away from the intersection center. It is approaching a crosswalk 4 meters ahead.
Similar to the aggregated context in Appendix A.5, the blue and orange sections contain answers to questions appearing in WOMD-Reasoning: blue addresses the Road Environment and Ego Agent Status, while orange covers Other Agent Status.
A.7 Ablation on Motion-LLaVA Training Strategy
The original LLaVA [28] training pipeline consists of two stages: a projector alignment stage (also known as pre-training) to align vision and language features, followed by a fine-tuning stage to enhance instruction-following capabilities. Throughout this process, the vision encoder remains frozen. Since our Motion-LLaVA builds on the architecture of the original LLaVA [28], it is intuitive to adopt a similar training approach for the WOMD-Reasoning task.
However, based on extensive experiments with fine-tuning strategies in Table 14, we find that incorporating projector alignment stage or keeping the motion vector encoder frozen throughout, which are practices employed in LLaVA [28], resulted in degraded performance on WOMD-Reasoning. Consequently, our Motion-LLaVA adopts a single-stage fine-tuning approach, where the motion vector encoder is unfrozen.